OpenStudy (anonymous):

What is the solution to the equation 9^(x + 1) = 27 ?

mathslover (mathslover):

\[\huge{9^{x+1}=27}\] \[\huge{3^{2(x+1)}=3^3}\] \[\huge{2(x+1)=3}\]

mathslover (mathslover):

can u solve further?

OpenStudy (anonymous):

2x+2=3 2x=1 x=1/2

OpenStudy (anonymous):

thank you!

mathslover (mathslover):

gr8 work @billybob3